Calculate the electric flux linked with a sphere of radius 1m and charge 1c at its centre

Grade:12

1 Answers

Khimraj
7 years ago
Electric flux through gaussian surface is given by ∅ = Q/ ε0 Where Q is charge enclosed by gaussian surface and ε0 = vacuum permittivity it`s valve is 8.85×10^–12 F/mSo flux = 1/ε0 = 1.13*10^11